a restaurant sells fries for RS 120 each.burger for RS 200 each and sandwitches for RS 250 each if the profit on fries is 1/3 of its selling price ,profit on burgers is RS 60 more than the profit on fries and profit on sandwiches is 125% of profit on burgers what will be the total profit for the sale of 5 fries 10 burgers and 15 sandwitches
Answer by greenestamps(13327)



profit on each fries: 1/3 of the selling price of RS 120 = RS 40

profit on each burger: RS 60 more than on fries = RS 40 + RS 60 = RS 100

profit on each sandwich: 125% of the profit on burgers = 1.25(RS 100) = RS 125

profit on 5 fries, 10 burgers, and 15 sandwiches: 5(40)+10(100)+15(125)=200+1000+1875 = 3075

ANSWER: RS 3075